A pendulum consisting of a uniform rod and concentrated mass at its free end is pivoted to the inertial frame through a revolPlease solve and show full working
Answer #1

1 M = 5 kg m = 3 kg L=1.2 m T = 100 N.m 0,= 60° - =300 2.

In the image , 0 initial position is marked as and final position as , = 60 Initially Finally the rotation is 300° .. o O2 =

Total mass moment of innertia I Ip t I c 8.64 kg-m Initially w,=0 .: Initial anguler kinetic energy O Exi The center of grav

At final state The angular kinetic energy / IX Wz = 4.32 WY Ekz J The potential energy of rod m xgx (Lea sing, = - 15.3 J [ r

W = 110 x 4.1897 = 4180) J ². Energy balance Ek + Ep, tW = Ekzt Ep, +(-66.27) of 66.27 +418.9= 4.32 wy 4.32 W 551.44 W2 11.3

Mx L x W 765.89 N .: Total centrifugal forc. F = 228.77 + 715.80 =Dg5.66 N As vector, the 2 forces are 270 Ø = 62- 270 30 Ws

. Reactim forces on the pivot FH = 497.83 N Ev=940.75 m If you have in the comments. any confusim with the solution, ask me